40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 227 of 339  Not logged in ELOG logo
ID Date Author Type Category Subject
  5680   Mon Oct 17 17:07:30 2011 steveUpdateSUSETMX oplev returning beam od 3 mm

 ETMX oplev had 6 mm diameter beam on the qpd.  I relayed the beam path with 2 lenses  to get back  3 mm beam on the qpd

BRC 037  -100 Bi _concave lens and PCX 25  200 VIS do the job. Unfortunately the concave lens has the AR 1064.

 

 

Attachment 1: ETMX-oplev.jpg
ETMX-oplev.jpg
  5679   Mon Oct 17 14:26:22 2011 MirkoUpdateCDSSeismic BLRMS channels, new RMS calculation

Quote:

[Rana, Koji, Mirko]

We looked into the CDS RMS block c-code as described in Rolfs RCG app guide. Seems the block uses a first order LP filter with a corner freq. / time of 20k execution cycles. There are also some weird thersholds at +-2000counts in there.

I was looking into implementing a hand-made RMS block, by squaring, filtering, rooting. The new RMS (left) seems nicer than the old one (bottom right). Signal was 141counts sinus at 4Hz.

Filters used: Before squaring: 4th order butterworth BP at given freq. & (new) 6th order inverse Chebyshew 20dB at 0.9*lower BP freq. and 1.1*upper BP freq. => about 1dB at BP freq.

                       After squaring: 4th order butterworth LP @ 1Hz.

C1PEM execution time increased from about 20us to about 45us.

Made a new medm screen with the respective filters in place of the empty C1PEM_OVERVIEW. Should go onto the sitemap.

New_RMS_vs_old_RMS.png

Original RMS LP is slower than 0.1Hz, see below for single LP at 0.1Hz in the new RMS. Original RMS is faster than single LP @ 0.01Hz

Original_RMS_LP_slower_than_0.1Hz.png

Some of the channels are recorded as 256Hz DAQ channels now. Need to figure out how to record these as 16Hz EPICS channls.

 Channels are now going into EPICS channels (e.g. C1:PEM-ACC1_RMS_1_3 ). Adapted the PEM_SLOW.ini file. Channels don't yet show up in dataviewer. Probably due to other C1PEM maschine

  5678   Mon Oct 17 11:40:44 2011 KojiUpdateLSCREFL165 removed from the table

REFL165 removed from the table for the C(V) test

  5677   Mon Oct 17 11:06:31 2011 MirkoUpdateCDSPiping data from c1lsc to c1oaf

Changed, recompiled, installed and restarted c1rfm and c1oaf to get the MC1-3 Pitch and Yaw data into the c1oaf model.
Also changed c1oaf to use MCL as a witness channel (as well as an actuator).
Added the changes to svn.

  5676   Mon Oct 17 10:43:14 2011 MirkoUpdateCDSCommited changes to c1rfm

I want to make changes to c1rfm. Found uncommited changes to it from Sept 27. Since we recompiled it since then it should be safe to commit them, so I did. See svn log for details.

  5675   Mon Oct 17 07:57:24 2011 steveUpdateSUSETMX sus damping restored
  5674   Sun Oct 16 05:35:18 2011 ranaUpdateComputer Scripts / ProgramsFailing to set SUS summary screen values

Quote:

Quote:

I am trying to run Rana's setSensors.py script, but am failing.  Any inspiration would be appreciated:

rosalba:SUS_SUMMARY>./setSensors.py 1001708529 500 .1 .25
['./setSensors.py', '1001708529', '500', '.1', '.25']
/cvs/cds/caltech/apps/linux64/python/lib64/python2.4/site-packages/nds/__init__.py:28: RuntimeWarning: No protocol specified, attempting protocol nds_v2
  super(daq, self).__init__(host, port)
Connecting NDS2 .... authenticate done
Traceback (most recent call last):
  File "./setSensors.py", line 81, in ?
    mean = acquire(x)
  File "./setSensors.py", line 73, in acquire
    daq.request_channel(chans[x])
Boost.Python.ArgumentError: Python argument types in
    daq.request_channel(daq, str)
did not match C++ signature:
    request_channel(_daq_t {lvalue}, daq_channel_t*)

I'm not exactly sure what the problem is.  Line 73, looks like it should have 2 arguments in the daq.request_channel, but even if I put in the "daq" variable (which is set a few lines above), I get the exact same error.  So...something else is wrong.  Ideas from someone who "speaks" python??

 My guess is that this has something to do with the NDS client version you're using.  Try running the script on a machine where pynds and nds-client are known to be compatible, like pianosa.

 Doesn't work on pianosa either. Has someone changed the python environment?

pianosa:SUS_SUMMARY 0> ./setSensors.py 1000123215 600 0.1 0.25
Traceback (most recent call last):
  File "./setSensors.py", line 2, in <module>
    import nds
ImportError: No module named nds

  5673   Sun Oct 16 02:30:00 2011 ranaUpdateElectronicsTesting REFL165

Unless the bias feedback circuit has been tuned for the 1 mm diode, its possible that you are seeing some C(V) effects. Its easy to check by looking at the phase response at 165 MHz v. the DC photocurrent. Then the feedback or feedforward gain can be tuned.

 

  5672   Sat Oct 15 17:06:20 2011 KojiUpdateLSCInstallation REFL165

REFL165 was installed on the AP table last night.

Meanwhile I found the DC power level at the REFL PDs were 0.8~1.2V if the PRM is aligned and the IFO is not locked.
This corresponds to 16~24mA (20~30mW). This is too big.

The HWP of the REFL path were adjusted so that we have 6~10mA (8~12mW) on each PDs.

  5671   Sat Oct 15 16:42:08 2011 KojiUpdateLSCTesting REFL165

Test results of new REFL165 (the first attachment)

- The resonant freq 166.2MHz, Q=57 (previous Q was ~7)

- If we believe the TF measurement, the transimpedance at the resonance is 7.8k [V/A] and the shotnoise intercept current of ~1mA.
The linearity of the peak was confirmed by changing the modulation level of the beam.

- There is a riddle: the white light test indicates 4.5k [V/A] and 2.8mA for those numbers.
There are big descrepancies from those by the TF measurements.


Further analysis of the descrepancies:

Using the noise measurements with different DC current levels, the transimpedance for each frequency can be reconstructed.

Does this indicate the satiration by the white light???

- The TF measurement shows consistent mag&phase relationship at the resonance (c.f. LISO fit).
So this steep resonance is not an artifact by a noise or glitch but the real structure of the electronics.

- The TF measurement has been done with the photocurrent of ~0.3mA, while the transimpedance measurement
with the white light illumination has the practical effect only when the DC photocurrent is larger than 1mA
because of the circuit noise. Does this higher photo current affected the resonance?

- The off-resonant transimpedance agree with the TF measurement as far as we can see with those measurements.
This may mean that the actual resonant structure has been affected in the white light measurement.
(i.e. not the saturation of the RF opamp which causes the change of the gain at any freq.)
Is the above mentioned higher DC current causing the change of the diode capacitance or other property of the diode or the inductors???

Attachment 1: REFL165_test_111014_KA.pdf
REFL165_test_111014_KA.pdf REFL165_test_111014_KA.pdf REFL165_test_111014_KA.pdf REFL165_test_111014_KA.pdf
Attachment 2: REFL165_transimpedance2.pdf
REFL165_transimpedance2.pdf
  5670   Sat Oct 15 16:01:26 2011 kiwamuUpdateIOOabout LOCKIN module

Quote from #5669

To make things faster, I think we can just make a LOCKIN which has 3 inputs: it would have one oscillator, but 6 mixers. Should be simple to make.

 I think the idea of having multiple inputs in a LOCKIN module is also good for the LSC sensing matrix measurement.

Because right now I am measuring the responses of multiple sensors one by one while exciting a particular DOF by one oscillator.

Moreover in the LSC case the number of sensors, which we have to measure, is enormous (e.g. REFL11I/Q, REFL33I/Q, REFL55I/Q, ... POY11I/Q,...) and indeed it has been a long-time measurement.

  5669   Sat Oct 15 10:58:32 2011 ranaUpdateIOOMC WFS Output Matrix determination

In order to save time and sanity, you should not measure the pitch ->yaw and yaw-> pitch. It makes things too complicated and so far is just not significant. In the past we do not use these for the matrix work.

i.e. there should just be a 3x3 pitch matrix and a 3x3 yaw matrix. Once the loops are working we could investigate these things, but its really a very fine tweak at the end. There are quite a few other, more significant effects to handle before then.

To make things faster, I think we can just make a LOCKIN which has 3 inputs: it would have one oscillator, but 6 mixers. Should be simple to make.

  5668   Sat Oct 15 04:53:41 2011 SureshUpdateIOOMC WFS Output Matrix determination

After we had a rough idea of what the output matrix looks like (see this elog),
I tried to measure the transfer function coefs (TFCs) between mirror degrees of freedom and the WFS sensors (WFS1, WFS2 and MC_Trans QPD)
I found that the TFCs that I obtained at 10.15 Hz did not have any resemblance to the previously identified output matrix.

The problem, I realised, arises because the various lockins used
in the C1IOO model do not have the same relative phase; So if we try to excite a mirror with one of them
and demodulate a sensor signal on any of the other lockins the resulting output would not have the correct phase
(relative to the 1st lockin output). As a result unless we can reset the phase of all the lockins
simultaneously, we cannot demodulate multiple signals at the same time. (Joe/Jamie, Is it possible to
reset/reinitialise the phase of the CLK signals of the lockings? )


To get around this problem Koji suggested that I use just one lockin and determine all the 36 elements of the transfer matrix with it one at a
time rather than six at a time. When I did that, I got results consistent with the previoulsly determined outmatrix. It, of course, takes six times longer.

The matrix I first got is this one

 

(Mag, Phase) WFS1P WFS2P MC_T_P WFS1Y WFS2Y MC_T_Y
MC1P 0.332 0.518 0.316 0.019 0.066 0.000
  5.832 1.892 8.180 38.285 8.807 0.000
             
MC2P 0.355 1.798 0.342 0.023 0.144 0.000
  72.977 76.683 76.804 -16.364 77.451 71.579
             
MC3P 0.352 0.394 0.254 0.036 0.023 0.000
  2.005 3.249 6.249 5.712 26.349 NAN
             
MC1Y 0.051 0.055 0.058 0.788 1.024 0.001
  15.979 -4.487 -9.707 2.642 1.276 0.000
             
MC2Y 0.142 0.044 0.130 1.966 0.579 0.017
  70.044 83.818 76.397 74.283 76.134 77.269
             
MC3Y 0.044 0.052 0.022 0.080 0.948 0.194
  22.932 14.227 -45.924 9.677 1.125 1.124
             
Which can be  recast as below          
             
             
Magnitude WFS1P WFS2P MC_T_P WFS1Y WFS2Y MC_T_Y
MC1P 0.332 0.518 0.316 0.02 0.07 0
MC2P 0.355 1.798 0.342 0.02 0.14 0
MC3P 0.352 0.394 0.254 0.04 0.02 0
MC1Y 0.05 0.05 0.06 0.788 1.024 0.001
MC2Y 0.14 0.04 0.13 1.966 0.579 0.017
MC3Y 0.04 0.05 0.02 0.080 0.948 0.194

Phase WFS1P WFS2P MC_T_P WFS1Y WFS2Y MC_T_Y
MC1P 5.8 1.9 8.2 38.3 8.8 0.0
MC2P 73.0 76.7 76.8 -16.4 77.5 71.6
MC3P 2.0 3.2 6.2 5.7 26.3 NA
MC1Y 16.0 -4.5 -9.7 2.6 1.3 0.0
MC2Y 70.0 83.8 76.4 74.3 76.1 77.3
MC3Y 22.9 14.2 -45.9 9.7 1.1 1.1

 

Note that when MC2 is excited all the sensors showed a response about 75 deg out of phase with the reference (MC1 --> WFS1_PIT ) This was traced to the fact that while there is a 28Hz Elliptic LP filter on

both MC1 and MC3, while it is absent on MC2.  The Transfer functions  below show the difference in the phase of their response

WFS1P_RespTo_MC1andMC2.pdf

 

Since the MC2 POS is used in servos involving MCL we cannot afford to install a 28 Hz LP filter into the MC2 coil drivers.  However a module with the 28 Hz ELP was switched on, in each of the

 MC2 PIT and YAW filter banks.   I then checked to see if this has affected the relative phase of variour sensors.  The Phase angle between I and Q on each sensor channel was checked and corrected. 

Below are the spectra with the "before" and "after" correction of phases.

Before:

 WFS1_IQphase20111015_1.pdf        WFS2_IQphase20111015_1.pdf

 

Obviously this needed adjustment to reduce Q phase.   

  After twealkng the angle "R":

WFS1_IQphase20111015_2.pdf      WFS2_IQphase20111015_2.pdf

 

And again determined the transfer matrix (below). 

( I , Q ) WFS1P WFS2P MC_T_P WFS1Y WFS2Y MC_T_Y
MC1P 0.236 -0.300 0.229 0.049 -0.008 0.000
  0.015 -0.004 -0.027 0.011 -0.019 0.000
             
MC2P -0.125 -0.962 -0.135 0.114 0.028 0.000
  0.007 -0.052 -0.028 -0.004 -0.002 0.000
             
MC3P -0.225 -0.254 -0.255 -0.026 -0.010 0.000
  0.004 -0.012 -0.010 0.009 0.002 0.000
             
MC1Y -0.059 -0.023 -0.040 0.460 0.705 0.001
  0.004 0.003 0.009 0.009 0.017 0.000
             
MC2Y 0.030 0.190 0.040 -1.144 -0.296 0.015
  0.007 0.006 -0.009 -0.038 -0.009 0.001
             
MC3Y 0.018 -0.108 -0.018 0.134 -0.832 -0.001
  0.017 0.005 0.001 0.006 -0.016 0.000
 

Magnitude WFS1P WFS2P MC_T_P WFS1Y WFS2Y MC_T_Y
MC1P 0.236 0.300 0.231 0.05 0.02 0
MC2P 0.125 0.964 0.138 0.11 0.03 0
MC3P 0.225 0.254 0.255 0.03 0.01 0
MC1Y 0.06 0.02 0.04 0.460 0.705 0.001
MC2Y 0.03 0.01 0.19 1.145 0.296 0.015
MC3Y 0.02 0.11 0.02 0.134 0.832 0.001

Phase WFS1P WFS2P MC_T_P WFS1Y WFS2Y MC_T_Y
MC1P 3.694 0.784 -6.778 13.1 66.67 #DIV/0!
MC2P -3.214 3.100 11.557 -2.05 -4.48 0
MC3P -1.020 2.665 2.158 -19.1 -10.76 NA
MC1Y -3.96 -6.45 -12.14 1.085 1.357 0.000
MC2Y 13.22 41.08 -2.6 1.887 1.706 4.987
MC3Y 42.69 -2.56 -3.73 2.652 1.068 0.000

 

This time the signals are all nearly in the same phase and in agreement with the  outmatrix estimate made earlier.

 

I plugged these TFCs into the matrix inversion code: wfsmatrix2.m.   And get the following inverse:

 

  WFS1P_Act WFS2P_Act MC_Trans_P_Act WFS1Y_Act WFS2Y_Act MC_TRANS_Y_Act
MC1P 1 -0.64        
MC2P -0.27 -1        
MC3P 0.98 -0.65        
MC1Y       -0.26 -1  
MC2Y       1 0.12  
MC3Y       0.16 0.07  

 

I have ignored the MC2_Trans_P and Y sensors for now.

  5667   Fri Oct 14 18:38:41 2011 kiwamuUpdateSUSC1:SUS-ETMX_SPDMon fixed

Quote from #5666

 Anyway, things are working now:

 Good job ! Thank you so much

  5666   Fri Oct 14 16:20:11 2011 ZachUpdateSUSC1:SUS-ETMX_SPDMon fixed

I offered to help Kiwamu with some of the 40m work. The first task was to figure out why the ETMX side OSEM monitor was so low, since we know that the depth is about right. It was showing ~0.13 V to the others' ~0.7 V.

TL,WR: There was a wire disconnected from the breakout panel on the side of the rack

I started by pulling the board out and checking to make sure that it was working properly. I injected a sine wave to the SIDE IN and found that it showed up in the signal coming out of the back (into the crate) just fine (see below). One strange thing I noticed while testing the board is that both inputs for each used channel of the MAX333 switches on the board are shorted to their respective outputs. That is, the switches seem to be open to BOTH 0 and 1 logic states. This seems counterintuitive, but perhaps there's something about how these work that I don't know.

board_works.png

 

Then I went about tracing the signal from the back of the crate to the breakout panel on the side of the rack. I opened it up, verified that the ribbon cable was transmitting correctly, and as I went to plug it back in I noticed that one of the wires---the correct one---had come completely undone.

rut_roh_raggie.png

The screw clamp appeared to be a bit finicky, as I had to loosen and tighten it a few times before it finally seemed to grab hold of the wire. It probably just wasn't tight in the first place and the wire was pulled out. Anyway, things are working now:

Screen_shot_2011-10-14_at_4.09.45_PM.png

  5665   Fri Oct 14 04:35:45 2011 kiwamuUpdateLSClocking tonight

The lock of DRMI wasn't stable enough to measure the sensing matrix. Failed.

PRMI and SRMI were okay and in fact they could stay locked robustly for a long time.

I added a new option in the C1IFO_CONFIGURE screen so that one can choose Signal-Recycled Michelson in carrier resonant condition.

Screen_shot_2011-10-14_at_4.24.01.png

 Additionally the orthogonalization of the I-Q signals on REFL55 should be done because it hasn't been done.

 

  5664   Thu Oct 13 23:58:38 2011 KojiUpdateLSCfixing REFL165

I already have reported in this entry that REFL165 shows too high DC output which does not depend on the light level on the diode.
Today I removed REFL165 from the table and inspected it.

The diode has been burnt as shown in the first picture (left).
The window is smoked, and the photo sensitive surface has been removed from its base. It moves in the can.

The burnt diode was replaced to the new one.
The new one shows ~30% better capacitance of ~50pF
and I had to increase the inductance from 14nH (i.e. 15nH//220nH) to 18nH.
After some struggles to increase/decrease the stray inductance by moving the SMD capacitors a little, the resonance is reasonably tuned to 166MHz.

The comprehensive test will be performed shortly.

Attachment 1: PA131612.jpg
PA131612.jpg
Attachment 2: PA131618.jpg
PA131618.jpg
  5663   Thu Oct 13 21:44:48 2011 MirkoUpdateCDSSeismic BLRMS channels, new RMS calculation

[Rana, Koji, Mirko]

We looked into the CDS RMS block c-code as described in Rolfs RCG app guide. Seems the block uses a first order LP filter with a corner freq. / time of 20k execution cycles. There are also some weird thersholds at +-2000counts in there.

I was looking into implementing a hand-made RMS block, by squaring, filtering, rooting. The new RMS (left) seems nicer than the old one (bottom right). Signal was 141counts sinus at 4Hz.

Filters used: Before squaring: 4th order butterworth BP at given freq. & (new) 6th order inverse Chebyshew 20dB at 0.9*lower BP freq. and 1.1*upper BP freq. => about 1dB at BP freq.

                       After squaring: 4th order butterworth LP @ 1Hz.

C1PEM execution time increased from about 20us to about 45us.

Made a new medm screen with the respective filters in place of the empty C1PEM_OVERVIEW. Should go onto the sitemap.

New_RMS_vs_old_RMS.png

Original RMS LP is slower than 0.1Hz, see below for single LP at 0.1Hz in the new RMS. Original RMS is faster than single LP @ 0.01Hz

Original_RMS_LP_slower_than_0.1Hz.png

Some of the channels are recorded as 256Hz DAQ channels now. Need to figure out how to record these as 16Hz EPICS channls.

  5662   Thu Oct 13 21:40:59 2011 ranaSummaryVACRecovery from the power shutdown is completed

 As it turns out Steve is not crazy in this particular instance: the vacuum computer, linux3 , has some issues. I can login just fine, but trying to open a terminal makes the CPU rail and the RAM max out and eventually the machine freezes.

Under KDE, I can open a terminal OK as root, but if I then try a 'su controls', the same issue happens. Let's wait for Jamie to fix this.

  5661   Thu Oct 13 20:25:32 2011 KatrinUpdateGreen LockingLPF transfer function YARM

It is a 4th order filter with cut of frequency of 120 kHz.

 

Design

designLPF.png

 

Measurement

 LC_LPF.TIF

  5660   Thu Oct 13 14:23:09 2011 steveUpdateSUSITMX oplev with 3 mm beam on qpd

 I replaced the JDSU-Uniphase 1125P by 1103P He/Ne laser. This new laser had 2.8 mW output yesterday. It degraded to 0.5 mW by this morning.

The beam size on the QPD is ~3 mm  This should give us  better sensitivity. These are not the perfect lenses at all, but we have them here.

On the other hand, there are still some coherence below 1 Hz, so the laser intensity noise or clipping dominating  this  part of the spectrum.

 

Attachment 1: ITMXoplev1103p#2.png
ITMXoplev1103p#2.png
Attachment 2: ITMXoplev.png
ITMXoplev.png
  5659   Thu Oct 13 03:22:53 2011 kiwamuUpdateLSCmeasurement of sensing matrix : just began

- status update on LSC activity :

The measurement of the LSC sensing matrix has begun. But no useful results yet.

 

 The measurement script (#4850) ran pretty well after I did some modifications to adopt the script to the latest LSC model.

However the SNR weren't so great particularly in REFL33 in the PRMI configuration.

So I will tune the amplitude of excitations and integration times tomorrow.

Currently the excitation is at 238.1 Hz, where no disturbing structures are found in the spectra.

  5658   Wed Oct 12 19:58:32 2011 KatrinUpdateGreen LockingPower splitter is unbalanced

The mini circuit power splitter ZFRSC-42S+ used at the YARM has no balanced output as it should have according to the data sheet.

@ 0.05MHz  the amplitude unbalance should be 0.03 dB

A quick measurement shows that there is a LO amplitude dependent unbalance:

LO amplitude input (Vpp)  unbalanced output (dB)
1.3 3.66
1.4 4.08
1.5 4.28
1.6 4.36

So my question is, shall I replace the power splitter just in case it is further degrading?

  5657   Wed Oct 12 18:54:02 2011 KatrinUpdateGreen Locking60 Hz oscillation due to broken BNC cable

There was a 60 Hz and 120 Hz oscillation on the green PDH photo diode output. After a long search, I could identify that

the source was a broken BNC cable which was connected to the photo diode. I exchanged that BNC cable and the 60 Hz

and 120 Hz are gone :-)

With the new cable the PD output was less noisy so that it was easier to achieve a better alignment of the light to the cavity.

The reflected power could be reduced from 40% to 30%. For perfect alignment the reflected power would be 20%.

  5656   Wed Oct 12 17:53:01 2011 kiwamuUpdateLSCBS actuator response : fitting done and histroy of delays

An update on calibration of the BS actuator : A fitting has been done.

(Fitting)

 I used LISO for fitting the complex transfer function.
Because the data points around 1 Hz didn't have big coherence a few data points, which had coherence of less than 0.9, were excluded.
Also the fitting of the Q-factor wasn't successful due to the lack of good data points around the resonance.So I left Q fixed to be 5 in the fitting.
 

(Fitting result)

G =  2.18060874008e-8 +/- 6.425e-10 (2.95%)
f0 =  1.0100491195 +/- 1.51e-2 (1.49%)  [Hz]
Q = 5 (fixed)
delay =  423.2753462089e-6 +/- 4.989e-6 (1.18%)  [sec]

bode_BSactuator.png

 

(History of delay)
 Because we have been observing several different amount of delays in different configurations, perhaps it is worth to summarize those numbers.
     description  delay [usec]   elog entry
       MICH lock (BS actuation)  423 this entry
       LSC feed forward path  127 #5218
       MICH lock (BS actuation)  600 #4638
      ALS on X arm (ETMX actuation)  330 #4196
      RFM (from c1lsc to c1sus)  125 #4153
      from ADC to DAC (all the front end machine)  38-110 #3961
      from ADC to DAC (c1sus)  124 #3838
     RFM (c1ioo and c1sus)   8-62 #3855

Quote from #5648
Tomorrow I will do a serious fitting.

  5655   Wed Oct 12 08:43:30 2011 steveUpdateSUSITMX oplev improved a bit

 Atm2 is before optical path adjustment. The idea was to remove possible clipping in vacuum.

Coherense significantly reduced below 4 Hz

Today I will replace the He/Ne laser 1125P with 1103P

 

Attachment 1: ITMXoplev.png
ITMXoplev.png
Attachment 2: ITMXoplevservo_ON.png
ITMXoplevservo_ON.png
  5654   Wed Oct 12 00:35:42 2011 kiwamuUpdateLSCTRY path realigned

The TRY (TRansmitted light from Y arm ) path was a bit realigned because there had been a small clipping.

This clipping was introducing offsets on the error signals of the C1ASS servo.

(Story)

During I was running the C1ASS servo on the Y arm I found every time after the auto-alignment is done there still remained a slight offset in the beam pointing,

I looked at the CCD camera which looks at the transmitted light and then introduced an intentional misalignment in ETMY in order to find an obvious clipping.

Indeed there was a clipping in horizontal direction. I checked through the optics on the Y end optical bench.

On the second mirror (beam splitter) the beam was on a very edge. So I steered the first steering mirror to fix it,

In addition to that an iris which is placed between the first and second mirror was also clipping the beam,

So I fully opened the aperture of the iris.

  5653   Tue Oct 11 21:23:51 2011 JenneUpdateLSCArm absl lengths

Quote:

[Katrin, Jenne]

We took the data for the new absolute length measurement of both arms, after the latest vent and move.  We will analyze soonly.  We had done a round of analysis,  but then Koji pointed out that our data wasn't so clean because the whitening filters were on (and saturated the ADC).  We now have the data (but not the analysis) for the better data with the WF off.

So our dirty-data preliminary number for the X arm is 37.73meters, which is 14cm different from our old length.  We were supposed to move by ~20cm, so....either this measurement is bad because the data sucked (which it did), or we are 6cm off.  Or both.

I'll do another analysis with the clean data for both arms later today/tomorrow.

After analyzing the cleaner data, I get the following:


Y_Length_long  =  37.757 meters

X_Length_long  =  37.772 meters

 

As stated in the wiki, the goal arm length was  L = 37.7974 m for each arm. 

So we're within 2cm for X, and within 4cm for Y.

According to Kiwamu's awesome tolerance calculation, we need to be within 2cm for each arm.  Given that we started out 20cm wrong for X and 25cm wrong for Y, we're a lot closer now, even though we aren't meeting our Yarm requirement yet.

Probably some Optickle action is in order, to see what these new lengths give us in terms of sideband phase and other stuff.

If you want more digits on my calculated numbers (which are probably meaningless, but I haven't done a careful error analysis), in my directory ...../users/jenne/Xarm and ..../users/jenne/Yarm run Xarm_find_peaks_and_length.m and Yarm_find_peaks_and_length.m  respectively.  These will output the lengths.

  5652   Tue Oct 11 19:11:25 2011 kiwamuUpdateLSCRe: BS actuator reponse at low frequency : measured

I think the precision due to the loop gain uncertainty is something like 0.1% at 0.1 Hz. It's not the issue.

The real issue was the loud motion of MICH, which degrades the coherence of the measurement.

Also last night I tried the fringe hopping technique and gave it up for several reasons.

 

(uncertainty due to the loop gain)

When MICH is locked, the signal at C1:LSC-MICH_OUT can be expressed in frequency domain by
 
     MICH_OUT = G / (1+G) * (1 / A) * X + G / (1+G) * (1 / H) * (1 / A) * S,                 [1]
 
where G is the open loop gain, A is the actuator response, H is the sensor transfer function (constant factor),
X is the natural (unsuppressed) motion of MICH and S is an excitation injected at C1:LSC-MICH_EXC.
When the natural motion of MICH X is smaller than the excited displacement S/H, dividing MICH_OUT by S gives
 
   [Transfer function] = S / MICH_OUT
                                 = (1+G) /G * H * A
 
At low frequency the open loop gain is always big, so that the transfer function can be approximated to
 
   [Transfer function] ~ H *A
 
This approximation is valid with a precision of 1/G.
In my case yesterday, the open loop gain at 0.1Hz was about 103 or more than that, so the uncertainty due to the loop gain was 0.1% or even less.
 

(Effect from the MICH motion)

In the equation [1], it is shown that the MICH motion X shows up together with the excitation signal.
Actually this MICH motion term was not completely negligible and eventually this term disturbs the measurement resulting in a low coherence.
In order to get a high coherence in the measurement, X should be smaller than the excited displacement S/H,
 
      X << S / H
 
This the reason why I had to inject a big excitation signal. Although the coherence around 1Hz turned out to be still low due to the loud natural motion in MICH.
The excitation was already close to 0.1 um level in terms of peak-to-peak displacement, and I wasn't able to increase it any more because the MICH signal would run into a nonlinear regime.
In the worst case I lost the lock due to a too much excitation.
 
 
(Fringe hopping technique)
 
Actually I tried and gave up this technique. That's why I did the in-loop measurement.
My feeling is that this technique is not suitable for the 40m.
What I tried was to flip the sign of the MICH control such that the fringe hops from the dark fringe to the neighbor bright fringe or vice versa.
Difference in the control signal (C1:LSC-MICH_OUT) was supposed to give us the amount of signal which drives the actuator by exactly quarter of the laser wave length.
However this technique turned out to be not good because
    (1) BS actuator is too strong
          => expected difference in the control signal is quite small.
          => \lambda / 4 / A ~ 12 counts, where A is the actuator DC response of about 2.2e-8 [m/counts].
   (2) MICH motion was too loud
         => I saw such a tiny 12 counts difference in the control signal, but once the hopping is done the control signal immediately fluctuated and it was really hard to precisely measure it.
         => It's simply because MICH was loud, and the actuator tried to suppress the motion and it resulted such an immediate signal fluctuation in the control signal

Quote from #5649

This seems like an error prone method for DC responses due to the loop gain uncertainty. Better may be to use the fringe hopping method (c.f. Luca Matone) or the fringe counting method

 

  5651   Tue Oct 11 17:32:05 2011 jamieHowToEnvironment40m google maps link

Here's another useful link:

http://maps.google.com/maps?q=34.13928,-118.123756

  5650   Tue Oct 11 15:19:17 2011 ranaHowToEnvironment40m map

The Kinemetrics dudes are going to visit us @ 1:45 tomorrow (Wednesday) to check out our stacks, seismos, etc.

40mLabMap.png40mLabMap.jpg

I put these maps here on the elog since people are always getting lost trying to find the lab.

  5649   Tue Oct 11 15:14:50 2011 ranaUpdateLSCBS actuator reponse at low frequency : measured

Quote:

The response of the BS actuator in a low frequency regime has been measured.

 

This seems like an error prone method for DC responses due to the loop gain uncertainty. Better may be to use the fringe hopping method (c.f. Luca Matone) or the fringe counting method

  5648   Tue Oct 11 03:35:16 2011 kiwamuUpdateLSCBS actuator reponse at low frequency : measured

The response of the BS actuator in a low frequency regime has been measured.

After the measurement I did a coarse fit to see if the low frequency data agree with the high frequency response which I have measured two weeks ago (#5583)
So far it shows a good agreement with the high frequency data (see the plot below). Tomorrow I will do a serious fitting.
Once the calibration of BS is done, the low frequency responses of ITMs, PRM and SRM will be done by simply exciting BS and comparing them (maybe at a couple of frequency points around 0.1Hz).
BSactuator.png

(Measurement)

 + With free swinging MICH, the sensor (AS55_Q) was calibrated into counts/m.

     => The peak-peak counts was about 110 counts. So the sensor response is about 6.5x108 counts/m

 + Locked Michelson with AS55_Q and the signal was fedback to BS.

 + Set the UGF high enough so that the open loop gain below 10 Hz is greater than 1.

 + With DDT's swept sine measurement, C1:LSC-MICH_EXC was excited with a big amplitude of 40 counts.

 + Took a transfer function from C1:LSC-MICH_OUT to C1:LSC-MICH_EXC.

 + Calibrated the transfer function into m/counts by dividing it with the sensor response.

Quote from #5641
One possible reason is that my actuator responses are not so accurate below 1Hz.
I will measure the DC response of all the actuators and it will completely determine the shapes of the actuator responses except for the region around the resonance.

  5647   Tue Oct 11 00:59:55 2011 SureshUpdateIOOPreliminary locking of WFS loops

[Kiwamu, Koji, Suresh]

After correcting several errors in the WFS loops, we turned them on today and saw them working!

A while back (last week actually) I noticed that the WFS1 and WFS2  QPD segments were numbered in a different order but that their input matrices did not reflect this change.  As result the WFS pitch and yaw definitions were pretty much mixed up.  However even after clearing this up the signals still showed significant amount of cross couplings. 

This problem was finally traced to the relative phase between I and Q channels of the WFS segments.  Koji suggested that I check the relative phase between all the segments to be sure.  I then repeated the procedure that Valera and I followed in our earlier elog # 5321 , and found that the phases indeed required to be adjusted.  The excitation of MCL was at 6Hz, 100mVpp, as before.   The WFS response after this was much improved i.e.  the pitch yaw cross couplings were not visible when we misalign the MC with sliders in MC_ALIGN.  And the magnitude of the response also increased since the signal was transferred from the Q to I channels.  The the phases were tweaked by hand till Q< 1% of I.  However when I repeated this measurement an hour later (I wanted to save the plots) I found that the phases had changed by a few percent! 

Koji noticed that the MC_REFL camera image showed significant intensity fluctuations and advised that we try a higher frequency and lower amplitude to avoid nonlinear effects in the WFS and in the MCL to PSL lock.  So we repeated the process at 20Hz and 20mVpp, introduced at the IN2 of the MC_Servo.  The fig below shows the level to which we reduced the signal in Q.

WFS1_IQphase20111010.pdf    WFS2_IQphase20111010.pdf

We then checked the relative phase between various quadrants by looking at the time series in dataviewer.  WFS2 Seg4 phase had to be flipped to bring it into phase with all the rest. 

WFS_IQ_RelativePhase.png

 

After this I tried to see the WFS response to moving the MC1 and MC3 with the sliders and determined the following relations:

Pitch WFS1 WFS2
MC1 + -
MC2 - -
MC3 + +

 

Yaw WFS1 WFS2
MC1 + +
MC2 - -
MC3 + -

 

Disregarding the MC2 for now and assuming arbitrary gains of 1 for all elements we inverted these matrices inserted them into the WFS_servo_outmatrix.  We then found that the with a sign flip on all elements the loops were stable.  In the servo filters we had turned on only the filter modules 3 and 4.  There was no low frequency boost.   We gradually increased gain till we saw a significant suppression of the error signal at low frequencies as shown below.  There was also an associated suppression of Intensity noise at REFL_DC after a single bounce from PRM.

 WFS_error_signal_Oct10.pdf        WFS_reduction_intensity_noise_Oct10.png

 

To see if the locks can actually realign the MC if it were manually misaligned, we turned the loops off and misaligned MC by moving MC3 pitch by 0.05 (slider position), and then turned on the loops.  The locks were reengaged successfully and the MC regained alignment as seen on the StripTool below:

WFS_recentering_Oct10.png

 

We can now proceed with the fine tuning the servo filters and understand the system better:

Q1:    Does the WFS (I to Q) phase drift rapidly?  How can we prevent it?

Q2:   How is that we do not see any bounce or roll resonances on the WFS error signals?

Q3:  How do we include the MC2 QPD into the WFS Servo?

I will proceed with determination of the actual transfer coefs between the MC DoF and the WFS sensors. 

 

 

  5646   Mon Oct 10 18:53:04 2011 KatrinUpdateGreen LockingMirrors whose angle of incidence is not 45

The angle of incidence of light is for some mirrors on the YARM end table different from 45° even though the mirrors are coated for 45°.

The mirrors below are useful if there are plans to replace these mirrors by properly coated ones.

 

Mirror
Angle of incidence (degree)
1st 1" mirror right after laser* 10
2nd 1" mirror right after laser 35
1st 2" steering mirror to vacuum system 15
2nd 2" steering mirror to vacuum system 28

 

* This is the new mirror as decribed on http://nodus.ligo.caltech.edu:8080/40m/5623

 

  5645   Mon Oct 10 16:32:18 2011 steveUpdateSUSUL sensor of ETMY is recovered

 I lost UL osem voltage this morning when I was checking the actual connection at rack ETMY

This after noon I disconnected  the 64 pins IDE connector from satelite amp at the rack, and the two 25 pins Dsubs at this juction board.

UL OSEM recovered after reconnecting these three connectors.

Atm3, bad connection.........noisy UL

 

Attachment 1: ETMY_UL.png
ETMY_UL.png
Attachment 2: ETMY_OSEM_UL.png
ETMY_OSEM_UL.png
Attachment 3: noisyETMY_UL.png
noisyETMY_UL.png
  5644   Mon Oct 10 15:41:56 2011 KojiUpdatePSLPMC aligned

[Koji Suresh]

The steering mirrors for PMC were aligned. The transmission went up from 0.779 to 0.852.

  5643   Mon Oct 10 13:52:04 2011 kiwamuUpdateLSCRE: First attempt to estimate mode matching efficiency using interferometer

Quote from #5640

"^2"s are missing in the second equation, but the calculation results seem correct.

PRX and PRY have different mode matching because of the Michelson asymmetry.
Are individually estimated mode matching indicates any sign of reasonable mode mismatch?
(The difference can be very small because the asymmetry is not so big.)

- Thank you for the correction. The missing square operation has been added correctly on the last entry (#5639).

- As for the individual MM efficiency,
   I was assuming that the MM solutions are the same for PRX, PRY and the real PRC, so I haven't carefully checked differences between those cavities.
   However as you mentioned the difference in those cavities can be tiny due to the small 3 cm Schnupp asymmetry.
   Anyway I will briefly check it to make me sure.
  5642   Mon Oct 10 12:14:00 2011 MirkoUpdateComputer Scripts / ProgramsIMC simulations

[Mirko, Kiwamu]

I tried to answer two questions regarding the IMC:

1. What is the coupling of fluctuations in the SB freq. to SB transmitted power?
2. What (if any) is the influence of the IMC on the AM?

I ran into some weird things regarding the corresponding optickle simulations:
1. There seems to be some artifact at the beginning of every simulation sweep.
2. The position of features depends on the parameters of the sweep.

I mailed Matt asking if he sees some error in the simulations

opticke_xaxis.png

Attachment 2: DC_power.png
DC_power.png
Attachment 3: DC_power_B.png
DC_power_B.png
Attachment 4: IMC_simulation.zip
  5641   Mon Oct 10 10:14:43 2011 ranaUpdateLSClength fluctuations in SRCL

 How does it make sense that the motion at 0.1 Hz of PRC is 10x larger than MICH?

EDIT by KI:

 That's actually the point which I was wondering at. One possible reason is that my actuator responses are not so accurate below 1Hz.
I will measure the DC response of all the actuators and it will completely determine the shapes of the actuator responses except for the region around the resonance.
In the process of producing the plot I was assuming that all the actuator response have a 1 Hz resonance with Q of 5.
However in reality this assumption is not true because the resonant frequency is different in each actuator.
  5640   Mon Oct 10 00:01:26 2011 KojiUpdateLSCFirst attempt to estimate mode matching efficiency using interferometer

"^2"s are missing in the second equation, but the calculation results seem correct.

PRX and PRY have different mode matching because of the Michelson asymmetry.
Are individually estimated mode matching indicates any sign of reasonable mode mismatch?
(The difference can be very small because the asymmetry is not so big.)

  5639   Sun Oct 9 17:13:46 2011 kiwamuUpdateLSCFirst attempt to estimate mode matching efficiency using interferometer

The efficiency of the mode matching (MM) to PRC (Power-Recycling Cavity) has been estimated by using the interferometer.

The estimated MM efficiency is about 74 % when losses in the cavity are assumed to be zero.

 

(Motivation)

 There had been an issue that the recycling gain didn't go to the designed high value of about 42  (#5541).
One of the possibilities is a low efficiency in the MM to PRC (also see #5541).
Although the MM efficiency had been measured using a beam scan ( see a summary on the wiki) a long time ago, it haven't been verified.
Therefore the MM has to be reviewed by using the real interferometer.

(Measurement)

 The concept of this measurement is observe the amount of the reflected light from a power-recycled cavity and estimate the MM efficiency based on the measured reflectivities.
 Since using the real PRC (consisting of BS, ITMs and PRM) could be a too complicated system for this measurement,
simpler cavities, namely Power-Recycled ITMX and ITMY (PRX and PRY), were used to examine the MM efficiency.
 The measurement goes in the following order :
    (1) Measurement of the amount of the single-bounce reflection from PRM with BS and ITMs misaligned.
    (2) Lock PRX or PRY to carrier resonance.
    (3) Alignment of PRX/Y to maximize the intracavity power. This time ASDC was used as a monitor of the intracavity power.
    (4) Measurement of the amount of the reflected light when the cavity is in resonance. The value in REFLDC was averaged in 100 sec.
     => done by tdsavg 100 C1:LSC-REFLDC_OUT
The same measurement was performed for both PRX and PRY.
 
 - locking parameters -
  Sensor = REFL11_I
  Whitening gain = 10 (30 dB)
  PRCL_GAIN = 2
  UGF ~ 200 Hz

(Analysis)

In order to estimate the relation between the MM efficiency vs. the reflected light, two models are considered:
   (1) simple model => no loss and no sidebands
   (2) sideband-included model => no loss but sidebands are taken into the account of the reflection.
 
(1) In the simple model the reflectivity Prefl / Pin is expressed by
         [Reflectivity]  = Prefl / Pin = Z * Rcav +  (1- Z) * Rprm
 
where Z is MM efficiency and Rprm is the reflectivity of PRM
and Rcav is the reflectivity of PRX/Y when it's resonance and it is defined by
         Rcav = | rprm - ritm t2BS|2 / |1 -rprm ritm t2BS |2
 
Tprm = 5.75% and Titm = 1.4 % are assumed in all the calculations.
In the first equation the first term represents the mode matched light and hence it couples with the cavity through Rcav.
The second term is the non-mode-matched light and because they are not interacting with the cavity they will be simply reflected by PRM through Rprm.
 
(2) In reality two phase-modulated light (11 MHz and 55 MHz) will behave differently from the carrier.
For example when the carrier is in resonance the sidebands will be anti-resonance against the cavity.
So that the amount of REFLDC will be slightly bigger because of the reflection of the sidebands.
 
      Prefl = Z * Rcav * Pc + Z * Ranti * Ps +  (1- Z) * Rprm * (Pc + Ps)
 
where Pc and Ps are the power in the carrier light and the sidebands respectively.
And Ranti is the reflectivity of the anti-resonance PRX/Y, which can be obtained by replacing the minus sign by the plus sign in the equation of Rcav shown above.
It is assumed that the sum of the carrier power and sidebands power is the incident power Pin = Pc + Ps.
The power in the carrier and the sidebands were estimated based on the OSA measurement (#5519), so that
          Pc / Pin = |J0(0.14)|2 * |J0(0.17)|2 =  0.976
          Ps / Pin = 2 * |J1(0.14)|2 + 2 * |J1(0.17)|2 =  0.024
 

(Results)

Here are the measured values in REFLDC

 -- Measurement 1 : PRX
    Single bounce from PRM = 4802.27 counts
       ==> the incident power = 5095.25 counts
    Reflected light from PRX = 4433.88 counts
      ==> Reflectivity = 0.8702
 
-- Measurement 2 : PRY
    Single bounce from PRM = 4833.05 counts
       ==> the incident power = 5127.05 counts
    Reflected light from PRX = 4444.48 counts
      ==> Reflectivity = 0.86672
 
On average the reflectivity of power-recycled ITM cavity was 0.868 with a standard deviation of  0.001744.
Actually the standard deviation estimated here is not fair because the measurement was done by only twice,
but my intention was that I wanted to see how the error can affect the estimation of the MM efficiency.
Here is a plot comparing the model curves and the measured values with 5 sigma error box (5 times of measured standard deviation).

mm_reflection.png

It is shown that the mode matching efficiency is 73.7 % when the sideband-included model is considered.
With the 5 sigma deviation it can go from 65% to 82% but it is still low and lower than the beam scan measurement ( see a summary on the wiki). 

Anyways the estimated MM efficiency with the sidebands effect included and without loss effect is

        MM efficiency = 73.7 +/- 1.7 % (1 sigma error)  or +/- 8.7 % (5 sigma error)

  5638   Sat Oct 8 04:41:07 2011 kiwamuUpdateLSClength fluctuations in SRCL

For a comparison, the length fluctuation of Signal-Recycled ITMX (SRX) and ITMY (SRY) have been measured.

Roughly speaking the length motion of SRX and SRY are as loud as that of PRCL.

Some details about the measurement and data analysis can be found in the past elog entry (#5582).

In the process of converting the raw spectra to the calibrated displacements the SRM actuator was assumed to have a resonance at 1Hz with Q = 5.

length_noise.png

(Notes on SRX/Y locking)

     Sensor = REFL11_I
     Actuator = SRM
     Demod. phase = 40 deg
     SRCL_GAIN = 20
     UGF = 100 - 200 Hz
     Resonant condition = Carrier resonance
     Whitening gain = 0 dB
     ASDC = 360 counts

Quote from #5582

The MICH and PRCL motions have been measured in some different configurations.

      + PRCL is always noisier than MICH.

  5637   Sat Oct 8 00:44:42 2011 kiwamuUpdateLSCcalibration of SRM actuator

The AC response of the SRM actuator has been calibrated.

 actuators.png
(Summary of the calibration results)
     BS = 2.190e-08 / f2     [m/counts]
     ITMX  = 4.913e-09 / f2   [m/counts]
     ITMY  = 4.832e-09 / f2   [m/counts]
     PRM   = 2.022e-08 / f2  [m/counts]
     SRM   = 2.477e-08 / f2  [m/counts]    ( NEW ! ) 
 
(Measurement)
The same technique as I reported some times ago (#4721) were used.
The Signal-Recycled ITMY was locked for measuring the actuator response.
Since the ITMY actuator had been already calibrated, first the sensor was calibrated into [counts/m] by exciting the ITMY actuator and then calibrated the SRM actuator with swept sine measurement.
 
 - - notes to myself
   SRCL GAIN = 2.2
   Sensor = REFL11_I
   Demod. phase = 40 deg
   Resonant condition = Carrier resonant
   Gain in WF = 0 dB

Quote from #5583
The AC responses of the BS, ITMs and PRM actuators have been calibrated.

 

  5636   Fri Oct 7 23:23:05 2011 kiwamuUpdateSUSSRM oplev was oscillating

The SRM oplevs were found to be oscillating because of a small phase margin.

I reduced the gains to the half of them. The peak which Steve found today maybe due to this oscillation.

Quote from #5630

The SRM bounce peak 18.33 Hz. Suresh helped me to retune filter through Foton, but we failed to remove it. 

  5635   Fri Oct 7 22:48:26 2011 SureshUpdateIOOMC2 Trans QPD spot size and incident power decreased

After centering the spot on the MC2, I started to adjust the spot position on MC_TRANS_QPD to center the beam on it.  I noticed that the spot size was about 3 to 4mm dia. because the 200mm lens was too close to the QPD.  I moved it back and decreased the spot size to about 1mm and the sensitivity to spot position increased.  However, Koji noted that the QPD sectors were near saturation, so I put in a ND=0.3 filter to reduce the incident power on the QPD.

At optimal alignment the current QPD_SUM is around 25k to 26k counts (factor of 2 down). Eventually the gain of the QPD ckts have to be reduced to prevent saturation, for the moment this is temporary fix.

The MC_TRANS_SUM trigger for MC autolocker is working fine no further change was required.

  5634   Fri Oct 7 22:41:05 2011 KojiConfigurationGeneralScript backup fixed

Script backup regularly runs on op340m by crontab,
but the true backup were not taken since Oct 16, 2010
as the backup program was looking at the old script directory.
/cvs/cds/script/backupScripts.pl was modified to look at the new script directory.

OLD COMMAND:

$command = "cd /cvs/cds/caltech; /usr/sbin/tar cfX - $EXCLUDE_LIST scripts | bzip2 > $ARCHIVEDIR/scripts_$curdate.tar.bz2";

NEW COMMAND:

$command = "cd /opt/rtcds/caltech/c1; /usr/sbin/tar cfX - $EXCLUDE_LIST scripts | bzip2 > $ARCHIVEDIR/scripts_$curdate.tar.bz2";

  5633   Fri Oct 7 22:31:53 2011 kiwamuUpdateSUSnoisy ULSEN on ETMY

Yesterday Koji was claiming that the rms monitor of ULSEN on ETMY didn't well go down.

Indeed something bad is happening on ULSEN of ETMY.

I guess it could be a loose connection.

ETMY_OSEMs.png

(The unit of Y axis in the plot is not true. Don't believe it !)

  5632   Fri Oct 7 19:06:46 2011 SureshUpdateIOOMC spot positions: checked and corrected.

Koji and Kiwamu had adjusted the MC beam axis slightly such that we can couple the MC output into the Y-arm without exceeding the current range of adjustability on PZT1.  This changed the centering of beam spots on MC mirrors.  I checked the mc-decentering make sure we have not made too big a compromise.  And since we can move MC2 spot position while maintaining the current positions on MC1 and MC3 decentering, we can atleast eliminate the A2L coupling on that mirror.  I used the scripts in $scripts$/MC/moveMC2/ to adjust the MC2 spot position.

Spot positions in mm (MC1,2,3 pit MC1,2,3 yaw) before adjustment:
    1.4674   -0.3548    1.0199   -1.5519    1.9834   -1.5971

After correcting MC2:

    1.4528    0.1431    0.9958   -1.2147    0.3823   -2.0163

After correcting MC1:

    1.3745    0.0669    0.8899   -1.5269    0.0296   -1.7314

 

The spot positions on MC1 and MC3 are very nearly (+/- 0.06 mm) same as before, while the MC2  decentering has been reduced close to zero.

A slight adjustment of the PZTs may be required to reset the beam pointing.

  5631   Fri Oct 7 17:35:26 2011 KatrinUpdateGreen LockingPower on green YARM table

After all realignment is finished, here are the powers at several positions:

 

DSC_3496_power.JPG

ELOG V3.1.3-